# Invariant measures for some dissipative systems from the Jacobi last multiplier

### v2

@@ 1 Introduction
It is known that standard Hamiltonian dynamics is formulated on phase spaces that are symplectic manifolds, i.e., even-dimensional smooth manifolds equipped with a non-degenerate and closed two-form $\omega$ [1]. A powerful feature of this construction is the invariance of the volume-form $\omega^{n}\equiv\omega^{\wedge n}$ under a Hamiltonian phase flow. This result which is called Liouville’s theorem makes way for the formulation of classical statistical mechanics. The descriptions of dissipative systems are, however, quite different. For instance, some dissipative systems may be described by employing certain ‘generalized’ Hamiltonian frameworks such as by resorting to conformal Hamiltonian dynamics [2, 3] or by formulating Hamiltonian dynamics on contact or cosymplectic manifolds (see for example, [4]) [^1] It may be remarked that conformal Hamiltonian dynamics and contact Hamiltonian dynamics can also be used to describe certain dynamical equations relevant in biology and pattern formation [5]..
@@
Apart from linearly-damped systems with constant damping strength, another class of systems where the second-order equation of motion breaks time-reversal invariance are systems of the Liénard class, being described by the equation [^2] It may be noted that despite the presence of the damping term ‘$f(x)\dot{x}$’, the dynamics may not always be dissipative. For example, the nonlinear Liénard system $\ddot{x}+kx\dot{x}+\omega^{2}x+\frac{k^{2}}{9}x^{3}=0$ admits oscillatory (non-decaying) solutions [14].
@@
where $f(x)$ and $g(x)$ are suitable (usually smooth) real-valued functions of the real variable $x$. Here, the damping strength may depend on the variable $x$, thereby giving rise to the possibility of nonlinear dynamics as well as that of limit cycles [15]. The phase-space flows admit a non-trivial divergence unlike conservative Hamiltonian systems respecting traditional Liouville’s theorem. As may be expected, such systems cannot be described using the standard Hamiltonian approach although some such systems may be described by using nonstandard forms of Lagrangian and Hamiltonian functions [8, 10] provided $f(x)$ and $g(x)$ satisfy the so-called Cheillini integrability condition:
@@
The aim of this paper is to present expressions for invariant phase-space measures for some dissipative systems, focusing on three classes of systems – (a) systems that are described by conformal Hamiltonian dynamics, (b) systems that are described by contact Hamiltonian dynamics, and (c) systems of the Liénard class as defined above. For this purpose, we shall make use of the notion of Jacobi last multipliers [16]. In particular, we will re-derive the so-called ‘canonical’ invariant measure for contact Hamiltonian systems [11] by using the formalism of Jacobi last multipliers. Further, we shall describe a certain generalization of conformal Hamiltonian dynamics on ‘exact’ symplectic manifolds (specifically, cotangent bundles) which will allow us to discuss Liénard-type systems. In particular, it will be shown that when $f(x)$ and $g(x)$ satisfy the Cheillini integrability condition (1.2) (see for example, [8]), it is possible to find analytical expressions for phase-space measures which are invariant under the corresponding dynamics.
The paper is organized as follows. In the next section [Sec. (2)], we will recall some basic definitions regarding Jacobi last multipliers and the Hamiltonian formalism on symplectic and contact manifolds. Following this, in Sec. (3), we shall present invariant phase-space measures for both conformal Hamiltonian dynamics as well as contact Hamiltonian dynamics. Finally, in Sec. (4), we will present a generalized version of conformal Hamiltonian dynamics which can describe Liénard-type systems [Eq. (1.1)] and will also present the explicit expressions for invariant phase-space measures when the Cheillini integrability condition (1.2) is satisfied. We will conclude the paper with an extended discussion in Sec. (5).
@@ 2.1 Jacobi last multiplier
The last multiplier which was introduced by Jacobi in 1844 (see the classic text [16]) has turned out to be a useful tool in analytical mechanics. On one hand, the notion of Jacobi last multipliers allows one to deduce a conserved quantity (at least, locally) for a planar dynamical system or equivalently, for a dynamical system on an $m$-dimensional phase space if $(m-2)$ conserved quantities are already known – in that case, the system can be reduced to a family of planar ones [16, 17]. On the other hand, a last multiplier allows one to find the Lagrangians describing certain second-order differential equations [18, 19, 20, 21, 22, 23] (see also, the older work [24]), most notably, some from the Liénard class of systems [22, 25, 26]. More recently, it has found use in the context of non-holonomic Lagrangian mechanics [27], non-divergence-free vector fields [28], and also in the context of Lie symmetries [29, 30] and integrability [31].
Consider a dynamical system, i.e., a system of first-order equations which go as $\dot{x}_{i}=X_{i}(x_{1},x_{2},\cdots,x_{m})$, where $i=1,2,\cdots,m$. Here, $x_{i}$ could be thought of as being the (possibly local) coordinates in some region $U\subseteq\mathbb{R}^{m}$ and ${X_{i}}$ are real-valued and differentiable functions defined on $U$. The dynamical system is described by a first-order vector field that goes as
@@
Note that we have the definition $X(F)=\frac{dF}{dt}$, for a function $F$ defined on $U$. Let us consider the volume-form $\Omega=dx_{1}\wedge dx_{2}\wedge\cdots\wedge dx_{m}$. The divergence of $X$ can be defined from the Lie derivative as £$\pounds_{X}\Omega=({\rm div}X)\Omega$; if the divergence is zero, then the vector field is volume preserving. In general, ${\rm div}X\neq 0$. We can now define a Jacobi last multiplier (see [27, 28] for discussions on the definition and results presented below).
@@
Proof – Since £$\pounds_{MX}\Omega=0$, from Cartan’s formula for the Lie derivative [^3] Here, one makes use of Cartan’s ‘magic’ formula for the Lie derivative: £$\pounds_{X}\alpha=d(\iota_{X}\alpha)+\iota_{X}(d\alpha),$ (2.4) where $\alpha$ is a differential form and $X$ is a vector field., we have $d(\iota_{MX}\Omega)=\pounds_{X}(M\Omega)=0$.
@@ 2.2.1 Hamiltonian dynamics
£$\pounds_{X_{H}}\omega=d(\iota_{X_{H}}\omega)+\iota_{X_{H}}d\omega=0,$ (2.10)
@@ 2.2.2 Conformal Hamiltonian dynamics
£$\pounds_{{X}^{\gamma}_{H}}\omega=-\gamma\omega,$ (2.11)
@@
$\dot{q}^{i}=X_{H}(q^{i})=\frac{\partial H}{\partial p_{i}},\quad\quad\dot{p}_{i}=X_{H}(p_{i})=-\frac{\partial H}{\partial q^{i}}-\gamma p_{i}.$ (2.14)
@@
and from Cartan’s formula, this implies £$\pounds_{\Delta}\omega=\omega$. With this background, let us furnish a formal definition of a conformal Hamiltonian system.
@@ 2.3 Hamiltonian dynamics on contact manifolds
A contact manifold is a pair $(\mathcal{M}_{c},\eta)$, where $\mathcal{M}_{c}$ is a smooth manifold of (real) dimension $12n+$ 1 and $\eta$ is a one-form satisfying
@@
£$\pounds_{X_{h}}(\eta\wedge(d\eta)^{n})={\rm div}X_{h}(\eta\wedge(d\eta)^{n}),$ (2.24)
@@ 2.4 Constant-damping systems
$m\ddot{q}+m\gamma\dot{q}+V^{\prime}(q)=0,\quad\gamma>0,\quad V(q)\in C^{\infty}(\mathbb{R},\mathbb{R}),$ (2.26)
@@ 2.4.2 Description via contact vector field
Thus, the equations of motion are obtained as $\dot{s}=X_{h}(s)$, $\dot{q}=X_{h}(q)$, and $\dot{p}=X_{h}(p)$, which gives
@@ 3.1 Conformal Hamiltonian dynamics
[…]
@@
Remarks: It turns out that if $V^{\prime}(q)\neq 0$, then the integrand of (3.6) is not an exact differential and consequently, the expression (3.6) is of limited utility except for the free particle. However, for constant-damping systems in a non-trivial potential, invariant measure(s) can be found by considering either a contact-geometric description [Sec (3.2.3) below] or via a description in terms of generalized conformal vector fields as we will show in Sec. (4.3.1); in the latter approach, the potential function must be carefully chosen so as to satisfy the Cheillini integrability condition (1.2).
@@ 3.2.1 Level set $\{h^{-1}(0)\}$
[…]
@@ 4.1 Generalized conformal vector fields
£$\pounds_{Y}\omega=-K\omega,$ (4.1)
and then there exists a (locally defined) function $H$ in $T^{*}Q$ such that $\iota_{Y}\omega=dH-K\theta$. Such a vector field $Y$ will be denoted by $X^{K}_{H}$.
@@
Consider the cotangent bundle $T^{*}\mathbb{R}$ which is naturally equipped with the tautological one-form $\theta$ leading to the symplectic two-form $\omega=d\theta$. If $X^{K}_{H}$ be a generalized conformal vector field and if the function $K$ defined on $\mathbb{R}$ is differentiable, then the condition £$\pounds_{X^{K}_{H}}\omega=-K\omega$ is equivalent to $\iota_{X^{K}_{H}}\omega=dH-K\theta$ because $dK\wedge\theta=0$.
Proof – Consider the two-dimensional phase space $\pi:T^{*}\mathbb{R}\rightarrow\mathbb{R}$. If $q$ be a coordinate on $\mathbb{R}$ and $\pi:(q,p)\rightarrow q$ be the corresponding ‘induced’ fiber coordinates on the cotangent bundle, then naturally, $\theta=-pdq$ (implying $\omega=d\theta=dq\wedge dp$) from the coordinate definition of the tautological one-form on $T^{*}\mathbb{R}$. Thus, given a differentiable function $K(q)$ on $\mathbb{R}$, one has $dK\wedge\theta=0$. Consequently, from Cartan’s formula (2.4) for the Lie derivative, one can write £$\pounds_{X^{K}_{H}}\omega=d(\iota_{X^{K}_{H}}\omega)+\iota_{X^{K}_{H}}(d\omega)$, implying that $-K\omega=d(\iota_{X^{K}_{H}}\omega)$. The two sides of this equation agree if we put $\iota_{X^{K}_{H}}\omega=dH-K\theta$ because $dK\wedge\theta=0$.
@@ 4.2 Geometric description of Liénard systems
4.3 Cheillini integrability condition and invariant measures
@@
Remarks: The expressions (4.9) and (4.14) are true only when the condition (4.8) is true; (4.8) is equivalent to (1.2) which is termed the Cheillini integrability condition (see for example, [8, 10, 22, 23]). Notice that generally two values of $l$ are obtained from (4.8) and therefore the invariant measures are not unique.
@@ 4.3.1 Constant-damping systems in quadratic and constant potentials
If one considers the case of constant damping as a special case of the Liénard system, then putting $f(q)=K(q)=\gamma>0$ implies that the Cheillini integrability condition (4.8) demands
@@
This generally gives two values for $l$, whose reality imposes the condition $m\gamma^{2}\geq 8\alpha_{0}$. For simplicity, choosing $\alpha_{0}=m\gamma^{2}/8$ so that there is only one value of $l$, namely, $l=-1/2$, the expression (4.14) gives
@@
subject to $p\neq\frac{m\gamma q}{2}$. Notice that we cannot treat general potentials here as they will not satisfy the Cheillini integrability condition (4.8).
Constant potential: If $V^{\prime}(q)=0$, the Cheillini integrability condition (4.8) implies $l=0,-1$. Referring to equation (4.14), one must discard $l=0$ and take $l=-1$, giving $M\sim 1/p$. Thus, the result (4.9) now corresponds to the result (3.7) obtained earlier for the free particle with constant damping.

### v3

@@ 1 Introduction
It is known that standard Hamiltonian dynamics is formulated on phase spaces that are symplectic manifolds, i.e., even-dimensional smooth manifolds equipped with a non-degenerate and closed two-form $\omega$ [1]. A powerful feature of this construction is the invariance of the volume-form $\omega^{n}=\omega^{\wedge n}$ under a Hamiltonian phase flow. This result which is called Liouville’s theorem makes way for the formulation of classical statistical mechanics. The descriptions of dissipative systems are, however, quite different. For instance, some dissipative systems may be described by employing certain ‘generalized’ Hamiltonian frameworks such as by resorting to conformal Hamiltonian dynamics [2, 3] or by formulating Hamiltonian dynamics on contact or cosymplectic manifolds (see for example, [4]) [^1] It may be remarked that conformal Hamiltonian dynamics and contact Hamiltonian dynamics can also be used to describe certain dynamical equations relevant in biology and pattern formation [5]..
@@
Apart from linearly-damped systems with constant damping strength, another class of systems where the second-order equation of motion breaks time-reversal invariance are systems of the Liénard class, being described by the equation [^2] It may be noted that despite the presence of the damping term $f(x)\dot{x}$’, the dynamics may not always be dissipative. For example, the nonlinear Liénard system $\ddot{x}+kx\dot{x}+\omega^{2}x+\frac{k^{2}}{9}x^{3}=0$ admits oscillatory (non-decaying) solutions [14].
@@
where $f(x)$ and $g(x)$ are suitable (usually smooth) real-valued functions of the real variable $x$. Here, the damping strength may depend on the variable $x$, thereby giving rise to the possibility of nonlinear dynamics as well as that of limit cycles [15]. The phase-space flows admit a non-trivial divergence unlike conservative Hamiltonian systems respecting traditional Liouville’s theorem. As may be expected, such systems cannot be described using the standard Hamiltonian approach although some such systems may be described by using nonstandard forms of Lagrangian and Hamiltonian functions [8, 10] provided $f(x)$ and $g(x)$ satisfy the so-called Chiellini integrability condition:
@@
The aim of this paper is to present expressions for invariant phase-space measures for some dissipative systems, focusing on three classes of systems – (a) systems that are described by conformal Hamiltonian dynamics, (b) systems that are described by contact Hamiltonian dynamics, and (c) systems of the Liénard class as defined above. For this purpose, we shall make use of the notion of Jacobi last multipliers [16]. In particular, we will re-derive the so-called ‘canonical’ invariant measure for contact Hamiltonian systems [11] by using the formalism of Jacobi last multipliers. Further, we shall describe a certain generalization of conformal Hamiltonian dynamics on ‘exact’ symplectic manifolds (specifically, cotangent bundles) which will allow us to discuss Liénard-type systems. In particular, it will be shown that when $f(x)$ and $g(x)$ satisfy the Chiellini integrability condition (1.2) (see for example, [8]), it is possible to find analytical expressions for phase-space measures which are invariant under the corresponding dynamics.
The paper is organized as follows. In the next section [Sec. (2)], we will recall some basic definitions regarding Jacobi last multipliers and the Hamiltonian formalism on symplectic and contact manifolds. Following this, in Sec. (3), we shall present invariant phase-space measures for both conformal Hamiltonian dynamics as well as contact Hamiltonian dynamics. Finally, in Sec. (4), we will present a generalized version of conformal Hamiltonian dynamics which can describe Liénard-type systems [Eq. (1.1)] and will also present the explicit expressions for invariant phase-space measures when the Chiellini integrability condition (1.2) is satisfied. We will conclude the paper with an extended discussion in Sec. (5).
@@ 2.1 Jacobi last multiplier
The last multiplier which was introduced by Jacobi in 1844 (see the classic text [16]) has turned out to be a useful tool in analytical mechanics. On one hand, the notion of Jacobi last multipliers allows one to deduce a conserved quantity (at least, locally) for a planar dynamical system or equivalently, for a dynamical system on an $m$-dimensional phase space if (m2)(m-2) conserved quantities are already known – in that case, the system can be reduced to a family of planar ones [16, 17]. On the other hand, a last multiplier allows one to find the Lagrangians describing certain second-order differential equations [18, 19, 20, 21, 22, 23] (see also, the older work [24]), most notably, some from the Liénard class of systems [22, 25, 26]. More recently, it has found use in the context of non-holonomic Lagrangian mechanics [27], non-divergence-free vector fields [28], and also in the context of Lie symmetries [29, 30] and integrability [31].
Consider a dynamical system, i.e., a system of first-order equations which go as $\dot{x}_{i}=X_{i}(x_{1},x_{2},\cdots,x_{m})$, where $i=1,2,\cdots,m$. Here, $x_{i}$ could be thought of as being the (possibly local) coordinates in some region $U\subseteq\mathbb{R}^{m}$ and $X_{i}$} are real-valued and differentiable functions defined on $U$. The dynamical system is described by a vector field that goes as
@@
Note that we have the definition $X(F)=\frac{dF}{dt}$, for a function $F$ defined on $U$. Let us consider the volume-form $\Omega=dx_{1}\wedge dx_{2}\wedge\cdots\wedge dx_{m}$. The divergence of $X$ can be defined from the Lie derivative as $\pounds_{X}\Omega=({\rm div}X)\Omega$; if the divergence is zero, then the vector field is volume preserving. In general, ${\rm div}X\neq 0$. We can now define a Jacobi last multiplier (see [27, 28] for discussions on the definition and results presented below).
@@
Proof – Since $\pounds_{MX}\Omega=0$, from Cartan’s formula for the Lie derivative [^3] Here, one makes use of Cartan’s ‘magic’ formula for the Lie derivative: $\pounds_{X}\alpha=d(\iota_{X}\alpha)+\iota_{X}(d\alpha),$ (2.4) where $\alpha$ is a differential form and $X$ is a vector field., we have $d(\iota_{MX}\Omega)=\pounds_{X}(M\Omega)=0$.
@@ 2.2.1 Hamiltonian dynamics
$\pounds_{X_{H}}\omega=d(\iota_{X_{H}}\omega)+\iota_{X_{H}}d\omega=0,$ (2.10)
@@ 2.2.2 Conformal Hamiltonian dynamics
$\pounds_{{X}^{\gamma}_{H}}\omega=-\gamma\omega,$ (2.11)
@@
$\dot{q}^{i}=X^{\gamma}_{H}(q^{i})=\frac{\partial H}{\partial p_{i}},\quad\quad\dot{p}_{i}=X^{\gamma}_{H}(p_{i})=-\frac{\partial H}{\partial q^{i}}-\gamma p_{i}.$ (2.14)
@@
and from Cartan’s formula, this implies $\pounds_{\Delta}\omega=\omega$. With this background, let us furnish a formal definition of a conformal Hamiltonian system.
@@ 2.3 Hamiltonian dynamics on contact manifolds
A contact manifold is a pair $(\mathcal{M}_{c},\eta)$, where $\mathcal{M}_{c}$ is a smooth manifold of (real) dimension $2n+1$ and $\eta$ is a one-form satisfying
@@
$\pounds_{X_{h}}(\eta\wedge(d\eta)^{n})={\rm div}X_{h}(\eta\wedge(d\eta)^{n}),$ (2.24)
@@ 2.4 Constant-damping systems
$m\ddot{q}+m\gamma\dot{q}+V^{\prime}(q)=0,\quad\quad\gamma>0,\quad V(q)\in C^{\infty}(\mathbb{R},\mathbb{R}),$ (2.26)
@@ 2.4.2 Description via contact vector field
Thus, the equations of motion are obtained as $\dot{s}=X_{h}(s)$, $\dot{q}=X_{h}(q)$, and $\dot{p}=X_{h}(p)$, which give
@@ 3.1 Conformal Hamiltonian dynamics
$\Omega|_{\Delta(H)\neq 0}=\exp\bigg(-n\int[\Delta(H)]^{-1}dH\bigg)\omega^{n},$ (3.6)
@@
Remarks: It turns out that if $V^{\prime}(q)\neq 0$, then the integrand of (3.6) is not an exact differential and consequently, the expression (3.6) is of limited utility except for the free particle. However, for constant-damping systems in a non-trivial potential, invariant measure(s) can be found by considering either a contact-geometric description [Sec (3.2.3) below] or via a description in terms of generalized conformal vector fields as we will show in Sec. (4.3.1); in the latter approach, the potential function must be carefully chosen so as to satisfy the Chiellini integrability condition (1.2).
@@ 3.2.1 Level set $\{h^{-1}(0)\}$
In the region of the (contact) phase space where $h=0$, (2.23) implies that the flow of the contact vector field $X_{h}$ should be confined to within that region as $h$ is conserved when $h=0$. In other words, the level set $\{h^{-1}(0)\}$ is invariant under the contact Hamiltonian dynamics. Formally, a Jacobi last multiplier is the solution of equation (3.8), giving $M=\exp\big(\int(n+1)\xi(h)dt\big)$. However, owing to the fact that the level set $\{h^{-1}(0)\}$ is a lower-dimensional subspace of the contact manifold $(\mathcal{M}_{c},\eta)$, i.e., $\phi:\{h^{-1}(0)\}\hookrightarrow\mathcal{M}_{c}$, where $\phi$ is the inclusion map, we have $\phi^{*}(\eta\wedge(d\eta)^{n})=0$. In other words, $\eta\wedge(d\eta)^{n}$ is not a volume-form on the level set $\{h^{-1}(0)\}$. As a result, in what follows, we will restrict our attention to the region of the contact phase space given by $\mathcal{M}_{c}\setminus\{h^{-1}(0)\}$.
@@ 4.1 Generalized conformal vector fields
$\pounds_{Y}\omega=-K\omega,$ (4.1)
and then there exists a (locally defined) function $H$ in $T^{*}Q$ such that $\iota_{Y}\omega=dH-K\theta$ if $dK\wedge\theta=0$. Such a vector field $Y$ will be denoted by $X^{K}_{H}$.
@@
Consider the cotangent bundle $T^{*}\mathbb{R}$ which is naturally equipped with the tautological one-form $\theta$ leading to the symplectic two-form $\omega=d\theta$. If $X^{K}_{H}$ be a generalized conformal vector field and if the function $K$ defined on $\mathbb{R}$ is differentiable, then the condition $\pounds_{X^{K}_{H}}\omega=-K\omega$ is compatible with $\iota_{X^{K}_{H}}\omega=dH-K\theta$ because $dK\wedge\theta=0$.
Proof – Consider the two-dimensional phase space $\pi:T^{*}\mathbb{R}\rightarrow\mathbb{R}$. If $q$ be a coordinate on $\mathbb{R}$ and $\pi:(q,p)\rightarrow q$ be the corresponding ‘induced’ fiber coordinates on the cotangent bundle, then naturally, $\theta=-pdq$ (implying $\omega=d\theta=dq\wedge dp$) from the coordinate definition of the tautological one-form on $T^{*}\mathbb{R}$. Thus, given a differentiable function $K(q)$ on $\mathbb{R}$, one has $dK\wedge\theta=0$. Consequently, from Cartan’s formula (2.4) for the Lie derivative, one can write $\pounds_{X^{K}_{H}}\omega=d(\iota_{X^{K}_{H}}\omega)+\iota_{X^{K}_{H}}(d\omega)$, implying that $K\omega=d(\iota_{X^{K}_{H}}\omega)$. The two sides of this equation agree if we put $\iota_{X^{K}_{H}}\omega=dH-K\theta$ because $dK\wedge\theta=0$.
@@ 4.2 Geometric description of Liénard systems
4.3 Chiellini integrability condition and invariant measures
@@
Remarks: The expressions (4.9) and (4.14) are true only when the condition (4.8) is true; (4.8) is equivalent to (1.2) which is termed the Chiellini integrability condition (see for example, [8, 10, 22, 23]). Notice that generally two values of $l$ are obtained from (4.8) and therefore the invariant measures are not unique.
@@ 4.3.1 Constant-damping systems in quadratic and constant potentials
If one considers the case of constant damping as a special case of the Liénard system, then putting $f(q)=K(q)=\gamma>0$ implies that the Chiellini integrability condition (4.8) demands
@@
This generally gives two values for $l$, whose reality imposes the condition $m\gamma^{2}\geq 8\alpha_{0}$. For simplicity, choosing $\alpha_{0}=m\gamma^{2}/8$ so that there is only one value of $l$, namely, l=1/2l=-1/2, the expression (4.14) gives
@@
subject to $p\neq-\frac{m\gamma q}{2}$. Notice that we cannot treat general potentials here as they will not satisfy the Chiellini integrability condition (4.8).
Constant potential: If $V^{\prime}(q)=0$, the Chiellini integrability condition (4.8) implies l=0,1l=0,-1. Referring to equation (4.14), one must discard $l=0$ and take l=1l=-1, giving $M\sim 1/p$. Thus, the result (4.9) now corresponds to the result (3.7) obtained earlier for the free particle with constant damping.